OpenStudy (anonymous):

solve for r. 2r^2-2/3r=0 enter multiple answers by separating them by a comma.

OpenStudy (anonymous):

r = 0 , 1/3

OpenStudy (anonymous):

that answe was not correct

OpenStudy (anonymous):

maybe question is \[2r^2-\frac{2}{3r}=0\]?

OpenStudy (anonymous):

yes thats the correct question. i just dont know how to make the squared symbol on here

jhonyy9 (jhonyy9):

6r*3=2 --- 3r*3=1 --- r*3=1/3 ---

OpenStudy (anonymous):

can u make the answer more clear

jhonyy9 (jhonyy9):

2r*2-2/3r=0 --- 6r*3-2=0 ---